Please help me solve this word problem :
The dimensions of a rectangular garden were 3m by 10m. When both dimensions were increased by equal amounts , the area of the garden doubled. Find the new dimensions.

Answer by mananth(16949)

width = 3
length = 10
area =30 m^2
(3+x)(10+x)=60
30+3x+10x+x^2=60
x^2+13x+30-60=0
x^2+13x-30=0
x^2+15x-2x-30=0
x(x+15)-2(x+15)=0
(x+15)(x-2)=0
x=-15,x=2
x=2, Ignore negative value -15
new dimensions
length = 10+2=12 m
width =3+2=5m
..
Check
12*5=60 m^2